Let f(x)=1/(3x^(2)+1). Find the open intervals on which f is concave up (down). Then determine the x-coordinates of all inflection points of f.

Use interval notation if you are asked to find an interval or union of intervals. (This link opens instructions below)

f is concave up on the intervals

f is concave down on the intervals

The inflection points occur at x =
1
Expert's answer
2012-11-02T10:18:28-0400
find inflection points we must find where f''=0
f'=-6x/(3x^2+1)^2
f''=(6 (-1+9 x^2))/(1+3 x^2)^3
f"=0 9x^2=1
x=1/3 or x=-1/3
inflection points occur at x =1/3 and x=-1/3to define is function concave up or down we must checkf"<0 or f">0 when
x<-1/3 then f">0 then f concave down -infinity<x<-1/3-1/3<x<1/3 f"<0 then f concave up
x>1/3 f"<0 then f concave down